# DiffLight: A Partial Rewards Conditioned Diffusion Model for Traffic Signal Control with Missing Data

### v1

@@ Conclusion
In this work, we only consider two missing patterns: random missing and kriging missing. While in the real world, the missing pattern in the traffic network is more complex. Meanwhile, we just adopt SFM which is similar to k-nearest neighbor (KNN) to impute the traffic data for baselines. In addition, we set $\gamma=0$ in the return $R_{t}$ and only condition on partial rewards in DiffLight which could lead to the short-sightedness of agents. Future work could explore the influence on performance in more different missing patterns even mixed missing patterns, adopt more different imputation methods, and find out a more far-sighted method to control the traffic signals under data-missing scenarios.
@@ B.2 Structure of STFormer
To simplify the expression, we omit the embedding of diffusion timestep, trajectory timestep and rewards in Equation 15, 16, 17 and 18. In the implementation, the embedding of diffusion timestep and trajectory timestep are added to every input of CCA, SSA and TSA, and the embedding of rewards is added to every input of SSA and TSA.

### v2

@@ Conclusion
In this work, we only consider two missing patterns: random missing and kriging missing. While in the real world, the missing pattern in the traffic network is more complex. Meanwhile, we just adopt SFM which is similar to k-nearest neighbor (KNN) to impute the traffic data for baselines. In addition, our approach is conditioned on partial rewards instead of returns which could lead to the short-sightedness of agents. Future work could explore the influence on performance in more different missing patterns even mixed missing patterns, adopt more different imputation methods, and find out a more far-sighted method to control the traffic signals under data-missing scenarios.
@@ B.2 Structure of STFormer
To simplify the expression, we omit the embedding of diffusion timestep, trajectory timestep and rewards in Equation 16, 17, 18 and 19. In the implementation, the embedding of diffusion timestep and trajectory timestep are added to every input of CCA, SSA and TSA, and the embedding of rewards is added to every input of SSA and TSA.